OpenStudy (anonymous):

Solve the inequality 5 is greater than or equal to(6-5x)/3>0 and express it as an interval.

OpenStudy (anonymous):

\[5\ge\frac{6-5x}{3}>0\] For the interval I got \[[-\frac{9}{5},0)\]

OpenStudy (anonymous):

What I wanted to know is I had to change the signs twice once when I divided by negative 5 and another because 0 is greater than -9/5 and the inequality had to be reversed... so I just want to check if I did the right thing by switching my signs twice :).

OpenStudy (anonymous):

this is chemistry department, i propose you to put your question in math department to get your answer soon :)